OpenStudy (anonymous):

c/-4-8=-48 ( "c" over -4 )

OpenStudy (anonymous):

can you plz be a bit clear on this?

OpenStudy (anonymous):

couldn't get the equation

OpenStudy (anonymous):

is it (c/-4) - 8 = 48 ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thats the smart way to write it haha sorry

OpenStudy (anonymous):

ok then th eanswer is c=-224

OpenStudy (anonymous):

okay thanks

OpenStudy (anonymous):

ur welcome

OpenStudy (anonymous):

9=14+m/2 (m over 2)

OpenStudy (anonymous):

m = -10 is the answer

OpenStudy (anonymous):

that was quick

OpenStudy (anonymous):

thanks :)

OpenStudy (anonymous):

im gud at that

OpenStudy (anonymous):

haha no problem. Would you like another to test your quickness? :)

OpenStudy (anonymous):

sure :) if i can , i will

OpenStudy (anonymous):

go on..type

OpenStudy (anonymous):

getting there hold your zebra's

OpenStudy (anonymous):

5/3j - 6 = 94

OpenStudy (anonymous):

is it 5/(3j) - 6 =94 ?

OpenStudy (anonymous):

um

OpenStudy (anonymous):

sure?

OpenStudy (anonymous):

then j=60

OpenStudy (anonymous):

new record haha

OpenStudy (anonymous):

haha :)

OpenStudy (anonymous):

i have 3 more

OpenStudy (anonymous):

be quick...

OpenStudy (anonymous):

one by one

OpenStudy (anonymous):

-2k/7 = 36

OpenStudy (anonymous):

it i s-126

OpenStudy (anonymous):

Woo hoo! Next one: z/6 + 11 = -49 :)

OpenStudy (anonymous):

whoa guess i got a new fan , huh :) am i right?

OpenStudy (anonymous):

Maybe

OpenStudy (anonymous):

okay , now the answer is -360

OpenStudy (anonymous):

that was fast. We wasnt even talking bout that haha. Thats all the questions i have. Hopefully that'll bring my grade up.

OpenStudy (anonymous):

Thaaaaaank-Youu!:D ( yes, i'm a fan ! ) haha

OpenStudy (anonymous):

yeah...maybe that wud...i hpoe u wud try practising these...all the best for ur grades:) ...n yeah , i knew that...'beahibbitts' (sorry dunno ur name)

OpenStudy (anonymous):

alexx haha . and yeah it's alll gooood now.

OpenStudy (anonymous):

okay alex...do well and all the best :) ...any more questions?

OpenStudy (anonymous):

nope i think thats all, thanks again:)

OpenStudy (anonymous):

yeah yeah...ur welcome...i like to solve these kinda questions u know:)

OpenStudy (anonymous):

Super fast :D

OpenStudy (anonymous):

" super fast" ..was this a title for me?

OpenStudy (anonymous):

Sure was

OpenStudy (anonymous):

whoa ...then i probably need to answer all math questions for my fan huh...?

OpenStudy (anonymous):

totally..

OpenStudy (anonymous):

hmm..lets see the next time wen u have questions...

OpenStudy (anonymous):

ill talk to you no doubt hahaha

OpenStudy (anonymous):

okay..i wud be back after 2 days then..

OpenStudy (anonymous):

ill be at the beach i think..

OpenStudy (anonymous):

but we'll see =)

OpenStudy (anonymous):

he he alright then

OpenStudy (anonymous):

Well...

OpenStudy (anonymous):

See ya Super Fast :)

OpenStudy (anonymous):

yeah..bye alex...

OpenStudy (anonymous):

oh hey dont slow down any okay?

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

haha think about it

OpenStudy (anonymous):

bye

OpenStudy (anonymous):

yeah..thinker thinks :)